Jack is now 7 times as old as Jill. If 10 years from now, he will 3 years as old as Jill, how old is Jack now?
Answer by Maths68(1474)

Let
Present age of Jack = j
Present age of Jill = i

Jack is now 7 times as old as Jill.
j=7i.......(1)
If 10 years from now,
after 10 years Jack will be = j+10
after 10 years Jill will be = i+10
Then
he will 3 years as old as Jill, how old is Jack now?
j+10=(i+10)+3
j+10=i+10+3
j=i+10+3-10
j=i+3
Put the value of j from (1) to above equation
7i=i+3
7i-i=3
6i=3
6i/6=6/3
i=1/2
Put the value of i in (1)
j=7i
j=7 * 1/2
j=7/2
j=3.5

Present age of Jack = j = 3 and half years old
Present age of Jill = i = half year old.